I. Calculate the cost of each financing source for Keyboard, Inc. The tax rate is 40%; ?.liesti. Debt Preferred Stock Common Stock Twelve-year bond, 6.75% annual coupon, current price of S960.81 Dividend $5.50, current price of S63.75 Beta of 1.45, excess market return of 5.63%, risk-free rate of 6.05% Currently, there are 121 million shares of common stock outstanding at $48 per share and four million shares of preferred stock outstanding. In addition, there is currently $956 million of debt outstanding. Calculate the weighted average cost of capital using current market weights a. b. The company is targeting a capital structure of 25% debt, 5% preferred stock, and 70% common equity. Recalculate the cost of capital based on target weights. Compare your answer to part B.
Answer #1

PV of bond = FV/(1+r)t + \sum Ct/(1+r)t Coupon = 6.75% * 1000 = 675
960.81 = 1000/(1+r)12 +   \sumCT/(1+r)t
Cost of Debt Rd = 7.25%
Cost of Preferred stock Rps= Dividend/Current Price = 5.5/63.75 = 8.627%
Cost of Equity Re = Rf  +  \beta * (Rp - Rf) = 6.05% + 1.45 * (5.63%) = 14.2135%
a) Total value of Equity = 121*48 million = $5808 mn (E)
Total Value of Preferred stock = 4 *48 milliion = $192 mn (P)
Total value of Debt = $ 956 mn (D)
Total Value = $6956 mn (V)
Cost of Capital WACC = (E/V)*Re + (P/V)*Rps + (D/V) * Rd *(1-t)= 12.69%
b. Target capital structure
25% debt, 5% preferred stock, 70 % Equity
Cost of Capital WACC = (E/V)*Re + (P/V)*Rps + (D/V) * Rd* (1-t) = 11.46%
WACC of second part is lower as ratio of debt increase and debt is cheapest of prefered stock and equity